How much energy is required to change a 60-g ice cube from ice at -40°C to steam at 140°C? XMJ
Answer #1

Specific heat(C​​​​​​ice) = 2.108 J/g-°C, Latent heat of fusion of ice (L) = 334 J/g, Specific heat(C​​​​​water) = 4.184 J/g-°C, Latent heat of Vaporization(V) = 2260 J/g, Specific heat(C​​​​​​steam) = 1.996 J/g-°C

Heat energy required to heat ice from -40°C to 0°C (Q1):

Q1 = mC​​​​​ice​​∆T = (60 g)*(2.108 J/g-°C)*(0°C - (-40°C)) = 5059.2 J

Heat energy required to melt ice (Q2):

Q2 = mL = (60 g)*(334 J/g) = 20160 J

Heat energy required to heat water from 0°C to 100°C(Q3):

Q3 = mC​​​​​water∆T = (60 g)*(4.184 J/g-°C)*(100°C - 0°C) = 25104 J

Heat energy required to Vaporize water(Q4):

Q4 = mV = (60 g)*(2260 J/g) = 135600 J

Heat energy required to heat steam from 100°C to 140°C(Q5):

Q5 = mC​​​​​Steam∆T = (60 g)*(1.996 J/g-°C)*(140°C - 100°C) = 4790.4 J

Total heat required = Q1 + Q2 + Q3 + Q4 + Q5 = 19713.6 J = 0.19 MJ